How much heat does a 15-gram piece of aluminum metal release as it cools from 260°C to 35°C?


1
Expert's answer
2021-04-15T04:25:16-0400
"Q=cm(t2-t1)"

c (Al)=904 J/(kg•deg)


"Q=904*0.015*(35-260)=-3051 J"